a room has a width of 10cm and height of 5cm. calculate the following
area in cm
area in m
Answer #1

Area(in cm) = Height * Width = 10 * 5 = 50 cm

1 cm = 0.01 m

1 cm = 0.0001 m

Therefore,

Area(in m2) = 50 cm² = 50 * 0.0001 = 0.005 m2